# thanks: Contributed Equallythanks: Contributed Equally

### v1

@@ Abstract
Introduction.— Advances in the classical simulation of quantum systems are of utmost importance as they underpin the benchmarking of quantum algorithms [1, 2, 3, 4, 5, 6] and enable rigorous tests of nascent claims reporting quantum advantage [7, 8, 9, 10, 11]. The state-of-the-art quantum simulation methods primarily revolve around two seemingly independent paradigms: stabilizer methods that can efficiently simulate systems with high entanglement, but struggle with non-Clifford gates [12, 13, 14]; or alternatively tensor network-based methods that are well suited to simulations with large number of non-Clifford gates, but scale poorly with entanglement [15, 16, 5]. Therefore, naturally one may seek to amalgamate these methods in such a way as to harness the advantages of both. Indeed, there has been a significant push towards developing such hybrid simulation approaches  [17, 18, 19, 20, 21, 22, 23, 24, 25], with recent work on the Stabilizer Tensor Network (STN) protocol [25] presenting an efficient technique for the simulation of both highly-magical systems with low entanglement, and highly entangled systems with low magic, loosely defined by the number of non-Clifford operations in the simulation. Building on this, our work demonstrates a marked improvement in the simulability of highly entangled circuits with a large number of $T$-gates, achieved by incorporating a tool commonly used in quantum error correction, namely magic state injection [26, 27]. Furthermore, our work provides the first demonstrations of the STN protocol being utilised for benchmarking circuits of interest, namely $T$-doped Clifford circuits and the Hidden Bit Shift circuit.
@@
By introducing the use of magic state injection to STNs, we have developed the Magic state injection Augmented Stabilizer Tensor network (MAST) simulation framework. To highlight the computational capabilities of the MAST technique, we benchmark MAST and STN against two commonly used circuits, namely $T$-doped Clifford circuits [13, 19, 28, 29] and the Hidden Bit Shift circuit [30, 13, 7, 31, 32]. A remarkable finding from our work is that MAST is capable of offering a considerable improvement in simulation cost for some classes of circuits. In particular, for random $T$-doped $N$-qubit Clifford circuits, MAST is able to efficiently simulate circuits with up to $N$ $T$-gates in polynomial time. Beyond this regime, we see a sharp phase transition in simulation cost as the $T$-gate density is scaled. Even in this intermediate-depth case, circuits prepared with magic state injection continue to outperform both standard STN and conventional Matrix Product State (MPS) methods until a saturation depth is reached. For the Hidden Bit Shift circuit we find that MAST outperforms STN for relatively few multiply controlled $Z$ (CCZ) gates. Of note however, is that for both MAST and STN, the bond-dimension appears to scale sub-exponentially, which is in stark contrast to both standard MPS methods and analogous resource costs in stabilizer simulators [13, 7]. Overall, we find that combining STNs with magic state injection is a powerful tool for simulating systems with moderate non-stabilizerness yet high entanglement.
@@
where $\nu_{i}$ are the coefficients of the basis states, |$\ket{\phi}$ is a stabilizer state as per Appendix A and the $D_{\hat{\imath}}$ are operators defined with respect to the destabilizer tableau (see Appendix B).
In this protocol, Clifford operations (the gate set generated by $\{H,S,CNOT\}$ gates) correspond to an update of the basis that can be kept track of in a regular stabilizer tableau [33], whilst non-Clifford operations can be represented as a sum over elements of the stabilizer basis and correspond to a controlled rotation on the MPS. Expectation values for operators in the stabilizer basis are found using standard MPS methods, which is in general efficient for an MPS in canonical form [15].
@@
A more detailed outline of the STN protocol, including how Clifford, non-Clifford and projection operations are performed can be found in Appendix B.
[…]
Figure 1: Magic state injection consists of local operations preparing a magic state, then injection using Clifford gates. If the magic state is measured to be |$\ket{1}$, a correction gate must be applied. This figure shows the magic state gadget for applying a $T$-gate. In this case, the correction is a Clifford gate, but this is not the case in general.
Magic State Injected Stabilizer Tensor Networks.— Building on the observation above, we note that in an arbitrary quantum circuit, non-Clifford operations may be replaced by a state injection gadget, shown in Figure 1, which uses a pre-prepared magic state and a Clifford injection circuit to apply non-Clifford operations [34]. The protocol finishes with measuring the ancilla magic state qubit and possibly applying a correction, which is a Clifford correction operation for the $T$ ($\pi/4$ rotation around $Z$) gate. In a simulator it is possible to predetermine the outcome of this measurement and delay the actual projection until the end of the circuit. We improve STN with the addition of magic state injection, referred to as MAST. When using MAST, magic state preparation and simulation of the resulting Clifford circuit may be performed in polynomial time. The primary cost of a simulation is hence delayed until the final projection step.
@@
Indeed, we observe stark differences in simulation cost between the different techniques in this setting, finding that MAST significantly outperforms STN. As can be seen in Figure 2(b)-(c) we are able to simulate up to $N$ layers with bounded average bond dimension using MAST (Region A), in contrast to the exponentially scaling bond dimension displayed by STNs. In the intermediate-depth case of $N\lesssim t\lesssim 1.5N$ (Region B) we observe a “phase transition”, with the MAST bond dimension increasing exponentially in the number of additional layers. Despite the increase in computational resources required, MAST continues to significantly outperform STNs throughout Region B. Finally, for depths $t\gtrsim 1.5N$ (Region C), the MAST bond dimension reaches the maximum value of $2^{N/2}$.
@@
Now, as is well known, the resources required to simulate an $N$-qubit Clifford circuit doped with $t$ $T$-gates via the standard stabiliser framework scale as $\mathcal{O}({\rm poly}(N)\exp(t))$, and become intractably large when $t=\mathcal{O}(N)$ [12, 13]. More generally, it is strongly believed that the simulation of Clifford circuits with $\mathcal{O}(N)$ non-Clifford resources is generically intractable [40]. In this regime, for example, it has been shown that doped Clifford circuits form approximate unitary designs to a precision exponentially small in the system size, reproducing long-time signatures of quantum chaos [41]. In contrast, we see that MAST is consuming only bounded classical resources up until a critical threshold, which is exactly the expected onset of intractable quantum phenomena. We stress, however, that the MAST protocol does not admit efficient sampling from the probability distributions (in say, the computational basis) characterising these random circuits. While local expectation values may be efficiently computed, projection operations in general cannot. Our results are therefore commensurate with recent works [42, 11] that show the efficient simulability of Pauli expectation values from “locally-scrambling” families of circuits.
[…]
@@
Figure 3: (a) Hidden Bit Shift Circuit. $O_{g}$ is a random phase gate constructed from $\{Z,CZ,CCZ\}$. $Z(s)$ encodes the hidden bit shift, and the final measurement will result in $s$ with 100% probability. (b) Simulation cost of the hidden-bit shift circuit on 40 qubits. Two cases are simulated, where the CCZ gates are decomposed into four $T$-gates + two ancillas, as per [13], and where they are decomposed into seven $T$-gates with no extra ancillas. Details of these two decompositions can be found in Appendix C. While conventional STNs are dependent on the decomposition used, MAST is indifferent. (c) Here we simulate the hidden bit shift circuit with much larger numbers of qubits and $T$-gates. The number of Clifford operations in $O_{g}$ is scaled proportionately with $N$ to ensure sufficient mixing. The 4 $T$ decomposition is used for these results. As in Figure 2, increasing the number of qubits for a fixed number of $T$-gates reduces the simulation cost. (d) Larger simulations using the 7 $T$ decomposition. Simulations were run on Dual 2.45GHz AMD EPYC processors. Each point is an average over 10,000 shots.
@@
Finally, we consider the task of numerically simulating the trivial circuit $U^{\dagger}U\ket{0}$. We find that for these simulations the bond-dimension is in general significantly greater than $1$, indicating that MAST does not universally determine the lowest possible resource cost required for any circuit. We do, however, observe that the resource cost is path dependent where one is free is to choose the order in which ancilla qubits are projected. For the particular case of random circuits, projecting pairwise from the middle of the ancilla register out towards either end of that register, one finds that the bond-dimension of the MPS is at most $2$ throughout the simulation. This is as a result of each pair of projections generating two CNOT cascades as per Appendix B.4 that cancel each other out such that the result after the pair of projections is a product state. A priori one will likely not know the path which minimises the cost of a given simulation. It is an interesting open question to determine a method of finding such an (approximately) optimal projection path.
@@
Compared to Ref. [7], which simulates up to 64 $T$-gates on $40$ qubits using an extended stabilizer simulator that scales exponentially with the number of $T$-gates, Figure 3(b) shows that our technique is able to simulate 160 $T$-gates on $40$ qubits. Simulations of the Hidden Bit Shift circuit [13, 7] typically use a four $T$-gate decomposition of the $CCZ$ gate, as these techniques scale only with the number of non-Clifford operations. However, this requires the addition of two ancilla qubits. To explore the effect of this, we also simulated identical Hidden Bit Shift circuits with the standard 7 $T$-gate decomposition of the Toffoli gate [43]. Details of these can be found in Appendix C. With the $7$ $T$-gate decomposition, the simulation time for STN is lower compared to the $4$ $T$-gate decomposition, while it remains unchanged for MAST despite the circuit containing significantly more $T$-gates. These results are explained by the ancilla qubits and extra entangling gates present in the $4$ $T$ CCZ, which increase the simulation time for STN.
In Figure 3(c) and (d), we increase the size of the circuits being simulated. As per the results for random $T$-doped Clifford circuits, we find that increasing the number of qubits in the Hidden Bit Shift circuit does indeed reduce the bond-dimension under MAST, allowing large circuits of $4000$ qubits and $320$ $T$-gates to be simulated on a classical computer. Conversely, STN performs worse at higher qubit counts, and the constant factor advantage of requiring fewer ancillas is drowned out by the exponential cost of magic and entangling gates. Finally, we note that it has recently been shown theoretically that the hidden-shift circuit can be simulated in polynomial-time [44], although this has not been demonstrated in simulation.
Conclusion and Outlook.— Exploring the boundaries of the classical simulability of quantum devices is a fundamental task of quantum information theory that has gained increasing practical relevance in the current era of intermediate-scale experimentally-realised quantum computers. Powerful simulation algorithms serve to not only benchmark these devices, but to rigorously test nascent claims of quantum advantage [9, 10]. Indeed, and in light of this, recent years have seen remarkable progress in the classical simulation of both noisy [45, 42, 46, 47, 48] and noiseless [49, 11] quantum computers. Beyond their implied practical advantages, these breakthroughs yield further insight into the nature of quantum computation by identifying with increasing fineness the regimes which are “actually quantum”.
Here, we have extended this body of work by introducing the Magic state injection Augmented Stabilizer Tensor network (MAST) method, and shown it is capable of simulating large circuits that are both highly entangled and magical. We specifically demonstrate this for random $T$-doped Clifford circuits, performing simulations with $\mathcal{O}(N)$ $T$-gates for up to $N=200$ qubits with bounded STN bond dimension. Recently, a similar observation was made for the Clifford Augmented Matrix Product State (CAMPS) protocol [50], which is a different amalgamation of stabilizer and tensor network methods [17, 17, 23, 24] to the STNs considered in this work. Unlike the simulation method devised in this work, the CAMPS-based method requires an optimisation subroutine whereas our method does not.
To further benchmark the utility of MAST, we also considered the Hidden Bit Shift circuit. For this system, we have been able to perform simulations with a significant number of qubits and non-Clifford operations: $N=4000$ qubits with $320$ $T$-gates.
@@
In addition, there is room to further refine MAST and STNs as simulation techniques. As shown in this work, the simulation cost is highly dependent on the decomposition of expensive operations and it may be possible to further optimise these decompositions, similar to the work done optimising magic state decompositions for stabilizer simulators [7, 54]. It may also be possible to replace the tensor network simulator with another simulation method.
Finally, an interesting open question relating to the theoretical underpinning of STN and by extension MAST is how the resource cost of these simulation methods relates to the quantum state that they are simulating, similar to how the bond-dimension of a conventional MPS simulation relates to the entanglement of the state, or how stabilizer rank relates to the magic of the state. This would not only allow for a more direct comparison with other simulation methods where the resource cost is well established but it would also shed a light on precisely what makes certain quantum systems difficult to simulate and narrow down the regime where quantum computers have an advantage over classical computers.
@@
Data Availability: All data generated in this work is presented in figures. Sources and data can be provided upon reasonable request to the corresponding authors.
@@ B.1 Clifford Operations
where $\tilde{D}_{\hat{\imath}}=C^{-1}D_{\hat{\imath}}C$. Recall that Clifford conjugation is the update rule for a stabilizer tableau simulator [12], so updating the tableau of |$\ket{\phi}$ is sufficient to update $D_{\hat{\imath}}$ as well. Hence Clifford operations only update |$\ket{\phi}$, using the standard stabilizer update rules [38], and leave the MPS unchanged.
@@ B.4 Projection
Appendix C $CCZ$ Gate Decompositions

### v2

@@ Abstract
Introduction.— Advances in the classical simulation of quantum systems are of utmost importance as they underpin the benchmarking of quantum algorithms [1, 2, 3, 4, 5, 6] and enable rigorous tests of nascent claims reporting quantum advantage [7, 8, 9, 10, 11]. The state-of-the-art quantum simulation methods primarily revolve around two seemingly independent paradigms: stabilizer methods that can efficiently simulate systems with high entanglement, but struggle with non-Clifford gates [12, 13, 14, 15]; or alternatively tensor network-based methods that are well suited to simulations with large number of non-Clifford gates, but scale poorly with entanglement [16, 17, 5]. Therefore, naturally one may seek to amalgamate these methods in such a way as to harness the advantages of both. Indeed, there has been a significant push towards developing such hybrid simulation approaches  [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], with recent work on the Stabilizer Tensor Network (STN) protocol [28] presenting an efficient technique for the simulation of both highly-magical systems with low entanglement, and highly entangled systems with low magic, loosely defined by the number of non-Clifford operations in the simulation. Building on this, our work demonstrates a marked improvement in the simulability of highly entangled circuits with a large number of $T$-gates, achieved by incorporating a tool commonly used in quantum error correction, namely magic state injection [29]. This is a natural next step for the STN formalism, as several simulation techniques [7, 30] have previously found an advantage by separating computation into special resource states and operations that can be simulated efficiently. Furthermore, our work provides the first demonstrations of the STN protocol being utilised for benchmarking circuits of interest, namely $T$-doped Clifford circuits and the Hidden Bit Shift circuit.
@@
By introducing the use of magic state injection to STNs, we have developed the Magic state injection Augmented Stabilizer Tensor network (MAST) simulation framework. To highlight the computational capabilities of the MAST technique, we benchmark MAST and STN against two commonly used circuits, namely $T$-doped Clifford circuits [13, 20, 31, 32, 15] and the Hidden Bit Shift circuit [33, 13, 7, 34, 35, 15]. A remarkable finding from our work is that MAST is capable of offering a considerable improvement in simulation cost for some classes of circuits. In particular, for random $T$-doped $N$-qubit Clifford circuits, MAST is able to efficiently simulate circuits with up to $N$ $T$-gates in polynomial time. Beyond this regime, we see a sharp phase transition in simulation cost as the $T$-gate density is scaled. Even in this intermediate-depth case, circuits prepared with magic state injection continue to outperform both standard STN and conventional Matrix Product State (MPS) methods until a saturation depth is reached. For the Hidden Bit Shift circuit we find that MAST outperforms STN for relatively few multiply controlled $Z$ (CCZ) gates. Of note however, is that for both MAST and STN, the bond-dimension appears to scale sub-exponentially, which is in stark contrast to both standard MPS methods and analogous resource costs in stabilizer simulators [13, 7]. Overall, we find that combining STNs with magic state injection is a powerful tool for simulating systems with moderate non-stabilizerness yet high entanglement.
@@
where $\nu_{i}$ are the coefficients of the basis states, |$\ket{\phi}$ is a stabilizer state as defined in Appendix A and the $D_{\hat{\imath}}$ are operators defined with respect to the destabilizer tableau.
In this protocol, Clifford operations (the gate set generated by $\{H,S,CNOT\}$ gates) correspond to an update of the basis that can be kept track of in a regular stabilizer tableau [36], whilst non-Clifford operations can be represented as a sum over elements of the stabilizer basis and correspond to a controlled rotation on the MPS. Expectation values for operators in the stabilizer basis are found using standard MPS methods, which is in general efficient for an MPS in canonical form [16]. Details on the implementation of STN can be found in the Appendix B.
@@
Figure 1: Magic state injection consists of local operations preparing a magic state, then injection using Clifford gates. If the magic state is measured to be |$\ket{1}$, a correction gate must be applied. This figure shows the magic state gadget for applying a $T$-gate (a $\pi/4$ rotation around $Z$). In this case, the correction is a Clifford gate, but this is not the case in general.
Magic State Injected Stabilizer Tensor Networks.— Building on the observation above, we note that in an arbitrary quantum circuit, non-Clifford operations may be replaced by a state injection gadget, shown in Figure 1. We improve STN with the addition of magic state injection, referred to as MAST. When using MAST, magic state preparation and simulation of the resulting Clifford circuit may be performed in polynomial time, noting that in simulation, it is possible to predetermine the outcome of the measurement in the state injection gadget and delay the actual projection until the end of the circuit. The primary cost of a simulation is hence delayed until the final projection step.
@@
In Appendix C we show that on average, MAST can simulate random $T$-doped Clifford circuits efficiently when $t$, the number of $T$ gates, is less than $N$, the number of qubits. The introduction of MAST is key to this argument, as it allows us to make assumptions about the structure of the constituent stabilizer tableau and tensor network in the simulator. Briefly, the argument is as follows; the cost of simulation in MAST comes from the projective measurement of magic states, which corresponds to an entangling operation in the MPS. We show that if this operation is on a certain subset of the MPS, it is not in fact entangling and does not increase the bond dimension. Finally, we show that for random $T$-doped Cliffords, there is a high probability of this being true when $t\lesssim N$.
@@
In Figure 2(b)-(c) we present numerical simulation results comparing MAST to STN, and show that we are able to simulate up to $N$ layers with bounded average bond dimension using MAST (Region A), in contrast to the exponentially scaling bond dimension displayed by STNs. In the intermediate-depth case of $N\lesssim t\lesssim 1.5N$ (Region B) we observe a “phase transition”, with the MAST bond dimension increasing exponentially in the number of additional layers. Despite the increase in computational resources required, MAST continues to significantly outperform STNs throughout Region B. Finally, for depths $t\gtrsim 1.5N$ (Region C), the MAST bond dimension reaches the maximum value of $2^{N/2}$.
@@
Now, as is well known, the resources required to simulate an $N$-qubit Clifford circuit doped with $t$ $T$-gates via the standard stabiliser framework scale as $\mathcal{O}({\rm poly}(N)\exp(t))$ and become intractably large when $t=\mathcal{O}(N)$ [12, 13]. More generally, it is strongly believed that the simulation of Clifford circuits with $\mathcal{O(N)}$ non-Clifford resources is generically intractable [42]. In this regime, for example, it has been shown that doped Clifford circuits form approximate unitary designs to a precision exponentially small in the system size, reproducing long-time signatures of quantum chaos [43]. In contrast, we see that MAST is consuming only bounded classical resources up until a critical threshold of $t\approx N$, which is exactly the expected onset of intractable quantum phenomena [42]. We also note that in the regime $t\leq N$, previous work has shown surprising advantages in unitary learning [44, 45] and entanglement manipulation [46]. Our results therefore adds to the growing list of useful properties of this ‘entanglement dominated’ phase. Noting that MAST allows for efficient computation of local expectations but does not admit efficient sampling from the probability distributions, our results appear commensurate with recent works [47, 11, 15] that show the efficient simulability of Pauli expectation values from “locally-scrambling” families of circuits. In particular, we note that while the stabilizer-based technique in [15] sees similar performance on this class of circuit and deterministic quantum circuits, it makes no improvement to the general problem of sampling from the probability distribution of a state, which scales as $\mathcal{O}(\exp(t)N^{3}w^{3})$, where $w$ is the number of bits being sampled (in the case of deterministic circuits we take $w=1$ for each qubit). In contrast, MAST scales as $\mathcal{O}(\exp(w))$ when $t<N$, so in the case where $t\gg w$, MAST will be more efficient at this computationally harder problem. In circuits where the final state has low entanglement, such as the Hidden Bit Shift below, sampling from the probability distribution will not increase the bond dimension and hence simulation cost, potentially giving MAST a large advantage over other techniques.
@@
Figure 3: (a) Hidden Bit Shift Circuit. $O_{g}$ is a random phase gate constructed from $\{Z,CZ,CCZ\}$. $Z(s)$ encodes the hidden bit shift, and the final measurement will result in $s$ with 100% probability. (b) Simulation cost of the Hidden Bit Shift circuit for up to $4000$ qubits. The number of Clifford operations in $O_{g}$ is scaled proportionately with $N$ to ensure sufficient mixing. In these simulations, the CCZ gate is decomposed into 4 $T$ gates and an ancilla as per Appendix D. Appendix E features results and discussion for an alternative decomposition of the CCZ gates. As in Figure 2, increasing the number of qubits for a fixed number of $T$-gates reduces the simulation cost.
@@
Finally, we consider the task of numerically simulating the trivial circuit $U^{\dagger}U\ket{0}$. We find that for these simulations the bond-dimension is in general significantly greater than $1$, indicating that MAST does not universally determine the lowest possible resource cost required for any circuit. We do, however, observe that the resource cost is path dependent where one is free to choose the order in which ancilla qubits are projected. For the particular case of random circuits, projecting pairwise from the middle of the ancilla register out towards either end of that register, one finds that the bond-dimension of the MPS is at most $2$ throughout the simulation. This is as a result of each pair of projections generating two CNOT cascades as per Appendix B.4 that cancel each other out such that the result after the pair of projections is a product state.
@@
Compared to Ref. [7], which simulates up to 64 $T$-gates on $40$ qubits using an extended stabilizer simulator that scales exponentially with the number of $T$-gates, Figure 3(b) shows that our technique is able to simulate 160 $T$-gates on $40$ qubits. In addition to this, we increase the size of the circuits being simulated. As per the results for random $T$-doped Clifford circuits, we find that increasing the number of qubits in the Hidden Bit Shift circuit does indeed reduce the bond-dimension under MAST, allowing large circuits of $4000$ qubits and $320$ $T$-gates to be simulated on a classical computer. Conversely, STN performs worse at higher qubit counts. Finally, we note that it has recently been shown theoretically that the hidden-shift circuit can be simulated in polynomial-time [48], although this has not been demonstrated in simulation.
Conclusion and Outlook.— The Magic state injection Augmented Stabilizer Tensor network (MAST) method is capable of simulating large circuits that are both highly entangled and magical. In particular, we highlight that random $T$-doped Clifford circuits can be efficiently simulated to depths less than the number of qubits and show numerical results with $\mathcal{O}(N)$ $T$-gates for up to $N=200$ qubits with bounded STN bond dimension. Recently, a similar observation was made for the Clifford Augmented Matrix Product State (CAMPS) protocol [49], which is a different amalgamation of stabilizer and tensor network methods [18, 18, 26, 27] to the STNs considered in this work. Unlike the simulation method devised in this work, the CAMPS-based method requires an optimisation subroutine whereas our method does not.
@@
In addition, there is room to further refine MAST and STNs as simulation techniques. As shown in this work, the simulation cost is highly dependent on the decomposition of expensive operations and it may be possible to further optimise these decompositions, similar to the work done optimising magic state decompositions for stabilizer simulators [7, 53]. This also leads to a theoretical understanding relating the simulation cost of MAST to a property of the quantum state being simulated, similar to bond-dimension and entanglement in a conventional tensor network, or stabilizer rank and magic. This would not only allow for a more direct comparison with other simulation methods where the resource cost is well established but it would also shed a light on precisely what makes certain quantum systems difficult to simulate and narrow down the regime where quantum computers have an advantage over classical computers.
@@
Data Availability: The implementation of the MAST protocol used to perform all numerical simulations featured in this work can be found in [54]. The data generated by the numerical simulations can be found in the figures.
@@ B.1 Clifford Operations
where $\tilde{D}_{\hat{\imath}}=CD_{\hat{\imath}}C^{-1}$. Recall that Clifford conjugation is the update rule for a stabilizer tableau simulator [12], so updating the tableau of |$\ket{\phi}$ is sufficient to update $D_{\hat{\imath}}$ as well. Hence Clifford operations only update |$\ket{\phi}$, using the standard stabilizer update rules [40], and leave the MPS unchanged.
@@ B.4 Projection
Appendix C Analysing the computational complexity of MAST
The cost of MAST simulation results from measurement projection, this can be of either the magic register or the data register. The tensor-network has the form |$\ket{\nu}=\ket{0}^{\otimes N}\prod_{i}^{t}R_{X}(\theta_{i})\ket{0}$. The projection operation for an operator $O=\alpha D_{\hat{a}}S_{\hat{b}}$ as per [28] is
$\frac{I+O}{2}$ $=\frac{1}{\sqrt{2}}\ket{0}\bra{0}_{k}(I\pm\alpha(-i)^{\hat{a}\cdot\hat{b}}X_{I_{X}}Y_{I_{Y}}Z_{I_{Z}})$
=|$=\ket{0}\bra{0}_{k}\mathcal{R}$ (15)
where $k$ is the first non-zero element of $\hat{a}$, $I_{Y}=\hat{a}+\hat{b}\mod{2}$, $I_{X}=\hat{a}+I_{Y}\mod{2}$ and $I_{Z}=\hat{b}+I_{Y}\mod{2}$. This operation can be rewritten as
$\frac{I+O}{2}$ $=\frac{1}{4\sqrt{2}}(I+Z_{k})\mathcal{R}$
$=\frac{1}{4\sqrt{2}}(\mathcal{R}+(Z_{k}\pm\alpha(-i)^{\hat{a}\cdot\hat{b}}Z_{k}X_{I_{X}}Y_{I_{Y}}Z_{I_{Z}}))$
$=\frac{1}{4\sqrt{2}}(\mathcal{R}+(Z_{k}\mp\alpha(-i)^{\hat{a}\cdot\hat{b}}X_{I_{X}}Y_{I_{Y}}Z_{I_{Z}}Z_{k}))$ (16)
where we have used that $k$ is either in $I_{X}$ or $I_{Y}$, but not both, and hence will anti-commute through $X_{I_{X}}Y_{I_{Y}}Z_{I_{Z}}$. There are two notable regions of interest here, one where $k<N$ and one where $k\geq N$. Applying Equation (16) to the former case results in
$\frac{I+O}{2}\ket{\nu}$ $=\frac{1}{4\sqrt{2}}(\mathcal{R}\ket{\nu}+(Z_{k}\ket{\nu}$
$\qquad\mp\alpha(-i)^{\hat{a}\cdot\hat{b}}X_{I_{X}}Y_{I_{Y}}Z_{I_{Z}}Z_{k})\ket{\nu})$
$=\frac{1}{4\sqrt{2}}(\mathcal{R}\ket{\nu}+(\ket{\nu}\mp\alpha(-i)^{\hat{a}\cdot\hat{b}}X_{I_{X}}Y_{I_{Y}}Z_{I_{Z}})\ket{\nu})$
$=\frac{1}{4\sqrt{2}}(\ket{\nu}\pm\alpha(-i)^{\hat{a}\cdot\hat{b}}X_{I_{X}}Y_{I_{Y}}Z_{I_{Z}}\ket{\nu}+\ket{\nu})$
$\qquad\mp\alpha(-i)^{\hat{a}\cdot\hat{b}}X_{I_{X}}Y_{I_{Y}}Z_{I_{Z}}\ket{\nu})$
$=\frac{1}{2\sqrt{2}}\ket{\nu}$
where we recognise that qubit $k$ is in the |$\ket{0}^{\otimes N}$ register and is hence unaffected by the $Z_{k}$ operation. In the regime where $k\geq N$ this simplification cannot be made in general and hence the projection operation may be entangling.
We can however recognise that in the $k>N$ case the only operations on the |$\ket{0}^{\otimes N}$ register are $Z_{I_{Y}}$, hence we can conclude that the data-register is completely unaffected by the projection operation which bounds the bond dimension by $2^{t/2}$, however one finds a significantly tighter bound by considering that there are only $N$ non-trivial stabilizer rows, and hence the logical projection operation occurs over at most $N$ qubits, one of which is projected in the tensor network, with any further logical projection occurring over a subset of the same $N$ qubits, except for the one that is projected, with an additional unentangled qubit now included. This means that at most there can be only $N$ entangled qubits after a logical projection, which bounds the bond dimension by $2^{N/2}$.
Given this, one may consider the computational complexity of MAST to be directly related to the location of the first destabilizer term in the projection operator. However, this is only an upper-bound, but as we will see this result is all that is needed to demonstrate the scaling observed for $T$-doped Clifford circuits seen in the main text.
C.1 The operator decomposition for T-doped Cliffords
[…]
Figure 4: The structure of the stabilizer tableau in MAST after executing a random circuit, but before projecting the register of magic state qubits. It consists of four regions, demarcated by the data and magic state registers. For our purposes, the right half of the tableau (i.e. the columns corresponding to the magic state qubits) are of interest. The lower right quadrant has diagonal $X$ terms, with $I$ terms on the off-diagonal. In the upper right quadrant, there is a random distribution of $X$ terms. The upper left quadrant of the table is a random tableau and not relevant to this discussion, while each row of the lower left quadrant is the identity Pauli string (i.e. $I^{\otimes N}$. This form may be found by observing that operations on the data register only modify the left half of the tableau, while magic state injection operations only add $X$ terms to the right half of the tableau.
22224444666688881010101012121212141414141616161618181818202020200.10.10.10.10.110.110.110.110.120.120.120.120.130.130.130.130.140.140.140.140.150.150.150.150.160.160.160.16Layers $(L)$$2\log(\chi)/N$$\sum_{i}^{L}2/2^{N-i}$
Figure 5: The expected bond-dimension for an $N=20$ $T$-doped Clifford circuit (in blue) along with the simulated bond-dimension using MAST (in orange). We find that the expected result is very close to the findings from the simulation. We note that our expression for the expected bond-dimension accounts only for a single increase in bond-dimension to $\chi=4$ as that is the primary mechanism by which bond-dimension increases for $t<N$.